# Information Structures in Stablecoin Markets

### v2

@@ Abstract
Stablecoins have historically depegged due from par to large sales, possibly of speculative nature, or poor reserve asset quality. Using a global game which addresses both concerns, we show that the selling pressure on stablecoin holders increases in the presence of a large redemption. While precise public knowledge reduces (increases) the probability of a run when fundamentals are strong (weak), interestingly, more precise private signals increase (reduce) the probability of a run when fundamentals are strong (weak), potentially explaining the stability of opaque stablecoins. The total run probability can be decomposed into components representing risks from large sales and poor collateral. By analyzing how these risk components vary with respect to information uncertainty and fundamentals, we can split the fundamental space into regions based on the type of risk a stablecoin issuer is more prone to. We suggest testable implications and connect our model’s implications to real-world applications, including depegging events, stablecoin regulations, and the no-questions-asked property of money.
@@ 2.2 Global Game
We assume that $\eta<1$ and $(1-\delta)\sigma_{x}<\sqrt{2\pi}\sigma^{2}$.
@@ 3 Results
$\mu>\bar{\theta}_{a}(x^{\star}_{a})+\frac{F^{-1}(1/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau}<0,$
@@
$\frac{\partial R_{1}}{\partial\tau}$$\frac{\partial R_{0}}{\partial\tau}$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)$-$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)+\frac{F^{-1}(1/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}}$$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)+\frac{F^{-1}(1/(1+\eta))}{2\tau_{x}\sqrt{\tau+\tau_{x}}}$
Figure 2: Diagram of how collateral risk and large sale risk change with respect to the precision of the common prior. The horizontal axis represents the prior mean $\mu$.
@@
$\mu>\bar{\theta}_{a}(x^{\star}_{a})+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau_{x}}>0,$
$\mu<\bar{\theta}_{a}(x^{\star}_{a})+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau_{x}}<0.$
@@
$\frac{\partial R_{1}}{\partial\tau_{x}}$$\frac{\partial R_{0}}{\partial\tau_{x}}$-$\overline{\theta}_{0}\left(x_{0}^{\star}\right)+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}$$\overline{\theta}_{1}\left(x_{1}^{\star}\right)+\frac{(2\tau+\tau_{x})\cdot F^{-1}(1/(1+\eta))}{2\tau\sqrt{\tau+\tau_{x}}}$++
Figure 4: Diagram of how collateral risk and large sale risk change with respect to the precision of private signals. The horizontal axis represents the prior mean $\mu$.

### v3

@@ Abstract
Stablecoins have historically depegged due from par to large sales, possibly of speculative nature, or poor reserve asset quality. Using a global game which addresses both concerns, we show that the selling pressure on stablecoin holders increases in the presence of a large redemption. While precise public knowledge reduces (increases) the probability of a run when fundamentals are strong (weak), interestingly, more precise private signals increase (reduce) the probability of a run when fundamentals are strong (weak), potentially explaining the stability of opaque stablecoins. The total run probability can be decomposed into components representing risks from large sales and poor collateral. By analyzing how these risk components vary with respect to information uncertainty and fundamentals, we can split the fundamental space into regions based on the type of risk a stablecoin issuer is more prone to. We suggest testable implications and connect our model’s implications to real-world applications, including depegging events, stablecoin regulations, and the no-questions-asked property of money. We are grateful for the support of Pantera Capital’s Catalyze Research Fellowship.
@@ 2.2 Global Game
We assume that $\eta<1$ and $(1-\delta)\sigma_{x}<\sqrt{2\pi}\sigma^{3}$.
@@ 3 Results
$\mu>\bar{\theta}_{a}(x^{\star}_{a})+\frac{2\tau+\tau_{x}}{3\tau_{x}\sqrt{\tau+\tau_{x}}}\cdot F^{-1}(1/(1+\eta))\implies\frac{dR_{a}}{d\tau}<0,$
@@
$\frac{\partial R_{1}}{\partial\tau}$$\frac{\partial R_{0}}{\partial\tau}$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)$-$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)+c$$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)+c$
Figure 2: Diagram of how collateral risk and large sale risk change with respect to the precision of the common prior. The horizontal axis represents the prior mean $\mu$, and $c=\frac{2\tau+\tau_{x}}{3\tau_{x}\sqrt{\tau+\tau_{x}}}\cdot F^{-1}(1/(1+\eta))\implies\frac{dR_{a}}{d\tau}<0$.
@@
$\mu>\bar{\theta}_{a}(x^{\star}_{a})+\frac{F^{-1}(1/(1+\eta))}{\sqrt{\tau+\tau_{x}}}\implies\frac{dR_{a}}{d\tau_{x}}>0,$
$\mu<\bar{\theta}_{a}(x^{\star}_{a})\implies\frac{dR_{a}}{d\tau_{x}}<0.$
@@
$\frac{\partial R_{1}}{\partial\tau}$$\frac{\partial R_{0}}{\partial\tau}$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)$$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)$$\overline{\theta}_{0}\!\left(x_{0}^{\star}\right)+c$$\overline{\theta}_{1}\!\left(x_{1}^{\star}\right)+c$
Figure 4: Diagram of how collateral risk and large sale risk change with respect to the precision of the common prior. The horizontal axis represents the prior mean $\mu$, and $c=\frac{F^{-1}(1/(1+\eta))}{\sqrt{\tau+\tau_{x}}}$.